# The local moduli of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres and invertible polynomials

### v1

@@ 2.2. Deformations of transverse holomorphic Sasakian structures
From these two theorems one can compute the dimension of the local moduli of Sasaki-Einstein metrics for certain links and therefore obtain lower bounds to the moduli space of Einstein metrics on many manifolds. We will do this for $\mathbb{Q}$-homology 7-spheres that are obtained as links of invertible polynomials. The key is to consider all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $d$ which is equivalent to find all the solutions of the following Diophantine equation
$w_{0}x_{0}+w_{1}x_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4}=d$
with variables $x_{i}\in\mathbb{Z}_{0}^{+}$ and where at least one of them is nonzero and subject to certain arithmetic constraints given by the type of invertible polynomials under study.
@@
3. Local moduli of Sasaki-Einstein metrics on $\mathbb{Q}$-homology 7-spheres
In this section we will compute the dimension of the local moduli of Sasaki-Einstein metrics of links that are $\mathbb{Q}$-homology 7-spheres. For that we will study the space of deformations of polynomials of cycle type, chain type, or certain iterated Thom-Sebastiani sums of these types. We only focus on these types of singularities for the reasons that were argued previously in Section 2.1: if a link $L_{f}$ is a $\mathbb{Q}$-homology 7-sphere and admits Sasaki-Einstein metric then $f$ has to be one of the aforementioned types. Since Sasaki-Einstein structures on the manifold determine Ricci-flat Kähler cone metrics on the corresponding affine cone, our results give information on the moduli of Calabi-Yau cones, see Problem 3.22 in [31] where the notion of K-moduli of Fano cones was introduced.
3.1. $\mathbb{Q}$-homology 7-spheres: the cycle type
@@
of degree $d$ and associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$, which is an element of Kollar’s list, such that $\gcd(d,w_{i})=1$ and $d=w_{0}+w_{1}+w_{2}+w_{3}+w_{4}-1$. From (3.1.1), we have the following relation between the weights $w_{i}$’s and the degree:
$a_{0}w_{0}+w_{4}=d,\ \ a_{1}w_{1}+w_{0}=d,\ \ a_{2}w_{2}+w_{1}=d,\ \ a_{3}w_{3}+w_{2}=d,\ \ a_{4}w_{4}+w_{3}=d$ (3.1.2)
Moreover, if the link $L_{f}$ is a $\mathbb{Q}$-homology sphere, it verifies that $d=1+a_{0}a_{1}a_{2}a_{3}a_{4}$. Using this, we can express each weight $w_{i}$ as:
$w_{i}=1-a_{i+1}+a_{i+1}a_{i+2}-a_{i+1}a_{i+2}a_{i+3}+a_{i+1}a_{i+2}a_{i+3}a_{i+4}$ (3.1.3)
where the subscript is taking mod $5$. Furthermore, considering the relations given in (3.1.2) and the fact that $\gcd(d,w_{i})=1$, we conclude that two consecutive weights $w_{i}$ and $w_{i+1}$ are always co-primes.
Now, we find the generators of the space of deformations of the orbifold $Z_{f}$. In [7], it was proven that the automorphism group of any orbifold $Z_{f}$ is finite as long as $w_{i}<\frac{1}{2}d$ for all but one of the $w_{i}^{\prime}s$ for $f$ quasi-smooth. Since the index $I=|\textbf{w}|-d=1$, the weight vector $\bf w$ does not admit polynomials that contain blocks of the form $z_{i}^{2}+z_{j}^{2}$. Thus for our $f$, $\dim\mathfrak{Aut}(Z_{f})=0,$ see [7]. Therefore, the complex dimension of $Z_{f}$ is given by
$\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}({d}))-\sum_{i}\dim H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i})).$ (3.1.4)
Let us compute $\dim H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}({d}))$. Here we consider all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $d$. Notice that it is equivalent to solve the following Diophantine equation
@@
First recall well-known result for Diophantine equations that will be used in the arguments that follow
@@
Let $\bf w$ defined as in (3.1.2). Then the Diophantine equation (3.1.5) has exactly five solutions. These solutions are
@@
where $q,r,s,t\in\mathbb{Z}$. Notice that to solve (3.1.5), we require $x_{i}\geqslant 0$ for all $i$. This implies that
@@
Replacing $a_{2}w_{2}=d-w_{1}$ in the above inequality, we obtain $-sw_{1}\leqslant w_{4}+td$. Moreover, as $d<2a_{1}w_{1}$ and $w_{4}=\min_{i}w_{i}\leqslant w_{1}$, we get
$-sw_{1}\leqslant w_{1}+td<w_{1}+2ta_{1}w_{1}\Rightarrow-s<1+2ta_{1}\Rightarrow-s\leqslant 2ta_{1}.$
@@
$-s\leqslant 2ta_{1}\leqslant 2a_{0}a_{1}.$ (3.1.16)
Also, from (3.1.11) we have $1+(t+s+r)a_{4}+qw_{3}\geqslant 0$, $-sa_{2}\geqslant r$ and $t\leqslant a_{0}$. Then
$-qw_{3}\leqslant 1+a_{4}(t+s+r)\leqslant 1+a_{4}(t-s(a_{2}-1))\leqslant 1+a_{4}(a_{0}-s(a_{2}-1)).$
@@
$-qw_{3}\leqslant 1+a_{4}(a_{0}+2a_{0}a_{1}(a_{2}-1))=1+a_{0}a_{4}-2a_{0}a_{1}a_{4}+2a_{0}a_{1}a_{2}a_{4}.$
Adding $01+a_{0}a_{4}-2a_{4}=1+a_{4}(a_{0}-2)>$ 0 to the right of the last inequality, we have
$-qw_{3}<2(1-a_{4}+a_{4}a_{0}-a_{4}a_{0}a_{1}+a_{4}a_{0}a_{1}a_{2})=2w_{3}\Rightarrow q>-2\Rightarrow q\geqslant-1.$
@@
    If $q=0$, we have the solution $(a_{0},0,0,0,1).$ Indeed from (3.1.11) we have
@@
    Now from $\gcd(w_{4},w_{0})=1$ and $w_{4}=\min_{i}w_{i}$, we have $w_{4}<w_{0}$. Then $-1<-\dfrac{w_{4}}{w_{0}}\leqslant t\leqslant a_{0}$. So we obtain $0\leqslant t\leqslant a_{0}$. In addition, from (3.1.17) we have $s\leqslant-ta_{1}\leqslant 0$. As $-t\leqslant s$, we have
@@
    If $q=-1$, we have to analyze for cases: In (3.1.11) we have
@@
    From (3.1.19), using $a_{0}-t\geqslant 0$ and $-sa_{2}\geqslant r$ in (3.1.20), we obtain
    $-sa_{2}+s\geqslant r+s\geqslant-1+a_{0}-t-a_{0}a_{1}+a_{0}a_{1}a_{2}\geqslant-1-a_{0}a_{1}+a_{0}a_{1}a_{2}.$ (3.1.21)
    Since $01-a_{2}<$ 0, then
@@
    Actually $s=1-a_{0}a_{1}.$ Indeed if $s=1-a_{0}a_{1}$ (which would force $a_{2}=2$) from (3.1.21) we can write
    $-s\geqslant r+s\geqslant-1+a_{0}-t+a_{0}a_{1}\geqslant-1+a_{0}a_{1}=-s.$
    In this inequality, we have that $t=a_{0}$. Thus, in the second inequality of (3.1.19) we obtain $s\leqslant-a_{0}a_{1}$, which contradicts the assumption $s=1-a_{0}a_{1}$. Therefore, we have $s=1-a_{0}a_{1}$ so we say
@@
      $-a_{0}w_{0}+a_{0}d+sd\geqslant-rw_{2}\geqslant sa_{2}w_{2}=sd-sw_{1}.$
      As $a_{1}w_{1}=d-w_{0}$, then $-a_{0}a_{1}\leqslant s$. Also, by (3.1.22) we know that $s\leqslant-a_{0}a_{1}$. Hence $s=-a_{0}a_{1}$. On the other hand, replacing $t=a_{0}-1$ and $s=-a_{0}a_{1}$ in (3.1.20), we obtain $r\geqslant a_{0}a_{1}a_{2}$. From (3.1.19) we have
@@
Next, we compute the generators of $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$. For this, we need to find all monomials $z_{0}^{a_{0}}z_{1}^{a_{1}}z_{2}^{a_{2}}z_{3}^{a_{3}}z_{4}^{a_{4}}$ of degree $w_{i}$. It is equivalent to solve the Diophantine equation
@@
Since $f$ is a cycle polynomial, we can work without loss of generality with $w_{i}=w_{1}$. Thus, the equation (3.1.25) results
@@
Remark 3.1.1.
@@
where $\alpha_{i}\in\mathbb{C}^{*}.$ The group $\mathfrak{G}_{\mathbf{w}}$ of complex automorphisms of $\mathbb{P}(\mathbf{w})$ is the projectivization of $G(\mathbf{w})$ which in this case is given by $\mathfrak{G}_{\mathbf{w}}=\left(\mathbb{C}^{*}\right)^{4}.$ Actually since the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the following quasi-smooth monomials $z_{0}^{a_{0}}z_{4},z_{0}z_{1}^{a_{1}},z_{1}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}}.$ It follows (c.f. [6]) that the moduli of the orbifold $Z_{f},$ a $\mathbb{Q}$-homology projective 3-space, is given by
@@
From the two previous lemmas, we obtain the following outcome in the context of Sasaki-Einstein structures for $\mathbb{Q}$-homology spheres.
Theorem 3.1.1.
Let $f$ be a cycle polynomial as in (3.1.1) of degree $d$ with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ such that $\gcd(d,w_{i})=1$ and index $I=|\textbf{w}|-d=1$. Then the complex dimension of the local moduli of the orbifold $Z_{f}=(f=0)/\mathbb{C}^{*}(\textbf{w})$, is equal to 00. Moreover, the generators of the spaces $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ and $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$ for all $i=0,\ldots,4$ are given in the previous lemmas of this subsection. Additionally, if $f$ belongs to one of the 236 $\mathbb{Q}$-homology spheres admitting Sasaki-Einstein metrics found in [5] and [13], then the dimension of the local moduli of Sasaki-Einstein metrics of $\mathbb{Q}$-homology 7-spheres at $L_{f}$ is zero dimensional. Thus $\mathbb{Q}$-homology 7-spheres given as links coming from polynomials $f$ as above do not admit inequivalent families of Sasaki-Einstein structures.
Proof.
The result follows from Lemmas 3.1.2, and 3.1.3. ∎
3.2. $\mathbb{Q}$-homology 7-spheres: Thom-Sebastiani sums of invertible polynomials

### v2

@@ 2.2. Deformations of transverse holomorphic Sasakian structures
From these two theorems it is clear that one can compute the dimension of the local moduli of Sasaki-Einstein metrics through the information given by the moduli of Kähler-Einstein metrics on Fano orbifolds. We will do this for rational homology 7-spheres that are obtained as links of invertible polynomials.
The procedure to determine the moduli of Kähler-Einstein metrics on Fano orbifolds goes as follows
  1. (a)
    Consider $X_{d}\subset\mathbb{P}\left(w_{0},\ldots,w_{n}\right)$ a well-formed and quasismooth weighted projective hypersurface of degree $d$ with finite automorphism group, cut out by an invertible polynomial $f$ of certain defined type.
  2. (b)
    Determine the linear system
    ${\mathcal{X}}_{d}=\left|\mathcal{O}_{\mathbb{P}(\bf w)}(d)\right|=\mathbb{P}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(d)),$
    that is, the parameter space for all hypersurfaces of degree $d$ in $\mathbb{P}(\mathbf{w}).$
  3. (c)
    Determine the automorphism group $\mathcal{G}(\bf w)$ of $\mathbb{P}(\mathbf{w}).$
  4. (d)
    Since $\mathcal{G}(\bf w)$ acts on $\left|\mathcal{O}_{\mathbb{P}(\bf w)}(d)\right|$ and there is an inclusion of the set $\mathcal{X}_{d}^{QS}$ of quasismooth hypersurfaces of degree $d$ in $\mathbb{P}(\bf w),$ as an open set in $\mathcal{X}_{d},$ one obtains the quotient
    $\mathcal{X}_{d}^{QS}/\mathcal{G}(\bf w),$
    which is a coarse moduli space, see [25] Corollary 1.2. Actually, since
    $\mathcal{X}_{d}^{QS}/\mathcal{G}({\bf{w}})\subset\mathbb{P}H^{0}(\mathbb{P}({\bf{w}}),\mathcal{O}(d))/\mathbb{P}G({\bf{w}})=H^{0}(\mathbb{P}({\bf{w}}),\mathcal{O}(d))/G({\bf{w}}),$
    where $G(\mathbf{w})$ is the group of automorphisms of the graded ring $S(\mathbf{w}),$ we will give a precise description of the former quotient.
  5. (e)
    Additionally, we will assume that the weighted hypersurface $X_{d}\subset\mathbb{P}\bf(w)$ satisfies the estimate $dI<\frac{n}{(n-1)}\min_{i,j}\left\{w_{i}w_{j}\right\}$ in Theorem 2.1.1, which implies the existence of Kähler-Einstein metrics in all the elements in $\mathcal{X}_{d}^{QS}.$ It follows from Theorems 2.2.1 and 2.2.2 that the number of parameters of the space of inequivalent Sasaki-Einstein metrics on the corresponding link $L_{f}=V_{f}\cap S^{2n+1}$ is given by the dimension of the moduli $\mathcal{X}_{d}^{QS}/\mathcal{G}(\bf w).$
Notice that well-formedness is required, otherwise we only obtain upper bounds in the dimension of the moduli space. For instance consider $Z_{9}\subset\mathbb{P}(3,3,3,3,3)$ a hypersurface of degree 9 which is isomorphic as a variety to $Z_{3}\subset\mathbb{P}(1,1,1,1,1)$ a hypersurface of degree 3 with the same equation. In Subsection 3.3, where we study the moduli of the Berglund-Hübsch duals, we deal with non well-formed weighted hypersurface where the procedure descried above suffices to determine the precise dimension of the moduli in most cases. In Section 4 we also discuss the role of the boundary points of $\overline{\mathcal{X}_{d}^{QS}/\mathcal{G}(\bf w)}$, that is, the non-quasismooth weighted varieties, in the construction of Sasaki-Einstein metrics on non-smooth links.
@@
3. Local moduli of Sasaki-Einstein metrics on smooth rational homology 7-spheres
In this section we will study the space of deformations of polynomials of cycle type, type I, type II and type III. We focus only on these types of singularities for the reasons that were argued in Section 2.1: if a link $L_{f}$ is a smooth rational homology 7-sphere and admits Sasaki-Einstein metric then $f$ has to be one of the aforementioned types. In the process we find the monomials generating $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(d))$ and the group of automorphisms of the weighted projective space $\mathbb{P}(\mathbf{w})$ that contains the corresponding orbifold.
3.1. Rational homology 7-spheres: the cycle type
@@
of degree $d$ and associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ with $\gcd(d,w_{i})=1,$ which in particular implies that the weighted hypersurface $Z_{f}\subset\mathbb{P}(\bf{w})$ is well-formed. From (3.1.1), we have the following relation between the weights $w_{i}$’s and the degree:
$a_{0}w_{0}+w_{4}=d,\ \ a_{1}w_{1}+w_{0}=d,\ \ a_{2}w_{2}+w_{1}=d,\ \ a_{3}w_{3}+w_{2}=d,\ \ a_{4}w_{4}+w_{3}=d.$ (3.1.2)
Moreover, if the link $L_{f}$ is a rational homology sphere, from the equality (2.1.1) $d=1+a_{0}a_{1}a_{2}a_{3}a_{4}$, we can express each weight $w_{i}$ as:
$w_{i}=1-a_{i+1}+a_{i+1}a_{i+2}-a_{i+1}a_{i+2}a_{i+3}+a_{i+1}a_{i+2}a_{i+3}a_{i+4},$ (3.1.3)
where the subscript is taken mod $5$. Furthermore, considering the relations given in (3.1.2) and the fact that $\gcd(d,w_{i})=1$, we conclude that two consecutive weights $w_{i}$ and $w_{i+1}$ are always co-prime.
Remark 3.1.1.
Notice that the assumption (e) below Theorem 2.2.2 and the relations in (3.1.2) imply that $a_{i}>1$ for $i=0,\ldots,4.$ This fact will be used throughout the proof of Lemma 3.1.2. Indeed, without loosing generality, let us assume that $a_{2}=1.$ Then the third equation in (3.1.2) gives $w_{1}+w_{2}=d$ and the estimate $dI<\frac{4}{3}\min_{i,j}\left\{w_{i}w_{j}\right\}$ can be written as
$(w_{1}+w_{2})(w_{0}+w_{3}+w_{4})<\frac{4}{3}\min_{i,j}\left\{w_{i}w_{j}\right\}.$
Since $6\min_{i,j}\left\{w_{i}w_{j}\right\}\leqslant(w_{0}+w_{3}+w_{4})(w_{1}+w_{2})$ we obtain a contradiction.
Now, we find the generators of the space of deformations of the orbifold $Z_{f}$. In [7], it was proven that the automorphism group of any orbifold $Z_{f}$ is finite as long as $w_{i}<\frac{1}{2}d$ for all but one of the $w_{i}^{\prime}s$ for $f$ quasismooth. Since $\gcd(d,w_{i})=1$, the weight vector $\bf w$ does not admit polynomials that contain blocks of the form $z_{i}^{2}+z_{j}^{2}$, thus $H^{0}(Z_{f},\Theta_{Z_{f}})=0.$ So the complex dimension of the moduli of $Z_{f}$ is given by
$\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}({d}))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i})).$ (3.1.4)
Let us compute $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}({d}))$. Here we consider all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $d$. Notice that it is equivalent to solving the following Diophantine equation
@@
First recall a well-known result for Diophantine equations that will be used in the arguments that follow.
@@
Let $\bf w$ be defined as in (3.1.2). Then the Diophantine equation (3.1.5) has exactly five solutions. These solutions are
@@
where $q,r,s,t\in\mathbb{Z}$. Notice that in order to solve (3.1.5), we require $x_{i}\geqslant 0$ for all $i$. This implies that
@@
Replacing $a_{2}w_{2}=d-w_{1}$ in the above inequality, we obtain $sw_{1}\leqslant w_{4}+td$. Moreover, as $d<2a_{1}w_{1}$ and $w_{4}=\min_{i}w_{i}\leqslant w_{1}$, we get
$sw_{1}\leqslant w_{1}+td<w_{1}+2ta_{1}w_{1}\Rightarrow-s<1+2ta_{1}\Rightarrow-s\leqslant 2ta_{1}.$
@@
$s\leqslant 2ta_{1}\leqslant 2a_{0}a_{1}.$ (3.1.16)
Also, from (3.1.11) we have $1+(t+s+r)a_{4}+qw_{3}\geqslant 0$, $sa_{2}\geqslant r$ and $t\leqslant a_{0}$. Then
$qw_{3}\leqslant 1+a_{4}(t+s+r)\leqslant 1+a_{4}(t-s(a_{2}-1))\leqslant 1+a_{4}(a_{0}-s(a_{2}-1)).$
@@
$qw_{3}\leqslant 1+a_{4}(a_{0}+2a_{0}a_{1}(a_{2}-1))=1+a_{0}a_{4}-2a_{0}a_{1}a_{4}+2a_{0}a_{1}a_{2}a_{4}.$
Adding $1+a_{0}a_{4}-2a_{4}=1+a_{4}(a_{0}-2)>0$ to the right of the last inequality, we have
$qw_{3}<2(1-a_{4}+a_{4}a_{0}-a_{4}a_{0}a_{1}+a_{4}a_{0}a_{1}a_{2})=2w_{3}\Rightarrow q>-2\Rightarrow q\geqslant-1.$
@@
    If $q=0$, we have the unique solution $(a_{0},0,0,0,1).$ Indeed from (3.1.11) we have
@@
    Now from $\gcd(w_{4},w_{0})=1$ and $w_{4}=\min_{i}w_{i}$, we have $w_{4}<w_{0}$. Then $1<-\dfrac{w_{4}}{w_{0}}\leqslant t\leqslant a_{0}$. So we obtain $0\leqslant t\leqslant a_{0}$. In addition, from (3.1.17) we have $s\leqslant-ta_{1}\leqslant 0$. As $t\leqslant s$, we have
@@
    If q=1q=-1, we have to analyze for cases: In (3.1.11) we have
@@
    From (3.1.19), using $a_{0}-t\geqslant 0$ and $sa_{2}\geqslant r$ in (3.1.20), we obtain
    $sa_{2}+s\geqslant r+s\geqslant-1+a_{0}-t-a_{0}a_{1}+a_{0}a_{1}a_{2}\geqslant-1-a_{0}a_{1}+a_{0}a_{1}a_{2}.$ (3.1.21)
    Since $1-a_{2}<0$, then
@@
    Actually $s<1-a_{0}a_{1}.$ Indeed if $s=1-a_{0}a_{1}$ (which would force $a_{2}=2$) from (3.1.21) we can write
    $s\geqslant r+s\geqslant-1+a_{0}-t+a_{0}a_{1}\geqslant-1+a_{0}a_{1}=-s.$
    In this inequality, we have that $t=a_{0}$. Thus, in the second inequality of (3.1.19) we obtain $s\leqslant-a_{0}a_{1}$, which contradicts the assumption $s=1-a_{0}a_{1}$. So we have $s<1-a_{0}a_{1}$ so we say
@@
      $a_{0}w_{0}+a_{0}d+sd\geqslant-rw_{2}\geqslant sa_{2}w_{2}=sd-sw_{1}.$
      As $a_{1}w_{1}=d-w_{0}$, then $a_{0}a_{1}\leqslant s$. Also, by (3.1.22) we know that $s\leqslant-a_{0}a_{1}$. Hence $s=-a_{0}a_{1}$. On the other hand, replacing $t=a_{0}-1$ and $s=-a_{0}a_{1}$ in (3.1.20), we obtain $r\geqslant a_{0}a_{1}a_{2}$. From (3.1.19) we have
@@
Next, we compute the generators of $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$. For this, we need to find all monomials $z_{0}^{a_{0}}z_{1}^{a_{1}}z_{2}^{a_{2}}z_{3}^{a_{3}}z_{4}^{a_{4}}$ of degree $w_{i}$. It is equivalent to solving the Diophantine equation
@@
Since $f$ is a cycle polynomial, we can work without loss of generality with $w_{i}=w_{1}$. Thus, the equation (3.1.25) results in
@@
Following [6], we collect the information given by the previous lemmas
Proposition 3.1.1.
@@
where $\alpha_{i}\in\mathbb{C}^{*}.$ The group $\mathcal{G}(\mathbf{w})$ of complex automorphisms of $\mathbb{P}(\mathbf{w})$ is the projectivization of $G(\mathbf{w})$ which in this case is given by $\mathcal{G}(\mathbf{w})=\left(\mathbb{C}^{*}\right)^{4}.$ Actually since the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the monomials $z_{0}^{a_{0}}z_{4},z_{0}z_{1}^{a_{1}},z_{1}z_{2}^{a_{2}},z_{2}z_{3}^{a_{3}},z_{3}z_{4}^{a_{4}}.$ Thus, the moduli of the orbifold $Z_{f}$ is included in the space
@@
$\square$
From the two previous lemmas, we obtain the following outcome in the context of Sasaki-Einstein structures for rational homology 7-spheres.
Proposition 3.1.2.
Let $f$ be a cycle polynomial as in (3.1.1) of degree $d$ with associated weight vector $\textbf{w}=(w_{0},w_{1},w_{2},w_{3},w_{4})$ such that $\gcd(d,w_{i})=1.$ Then the complex dimension of the moduli of the orbifold $Z_{f}=(f=0)/\mathbb{C}^{*}(\textbf{w})$, is equal to 0. Moreover, the generators of the spaces $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ and $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(w_{i}))$ for all $i=0,\ldots,4$ are given in Proposition 3.1.1. Additionally, if $f$ belongs to one of the 236 rational homology spheres admitting Sasaki-Einstein metrics found in [5] and [14], then the dimension of the local moduli of Sasaki-Einstein metrics of $\mathbb{Q}$-homology 7-spheres at $L_{f}$ is zero dimensional. Thus rational homology 7-spheres given as links coming from polynomials $f$ as above do not admit inequivalent families of Sasaki-Einstein structures.
$\square$
Example 3.1.1.
Consider the following cycle polynomial $f=z_{4}z_{0}^{2}+z_{0}z_{1}^{8}+z_{1}z_{2}^{4}+z_{2}z_{3}^{30}+z_{3}z_{4}^{3}$ which can be found in the Johnson and Kollár list of anticanonically embedded Fano Kähler-Einsten 3-folds. It follows that the corresponding weight vector is $\textbf{w}=(1945,477,1321,148,1871)$ and the degree is $d=5761$ so $\gcd(d,w_{i})=1$. As shown in [6] the corresponding link $L_{f}$ is a Sasaki-Einstein rational homology 7-sphere. By Proposition 3.1.1, the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the set of monomials $\{z_{4}z_{0}^{2},z_{0}z_{1}^{8},z_{1}z_{2}^{4},z_{2}z_{3}^{30},z_{3}z_{4}^{3}\}$. Furthermore, the moduli of the orbifold $Z_{f}$ is included in
$Span\{z_{4}z_{0}^{2},z_{0}z_{1}^{8},z_{1}z_{2}^{4},z_{2}z_{3}^{30},z_{3}z_{4}^{3}\}/G(\textbf{w})$
where $G(\textbf{w})$ is defined on generators by
$\varphi_{\mathrm{w}}\left(\begin{array}[]{l}z_{0}\\ z_{1}\\ z_{2}\\ z_{3}\\ z_{4}\end{array}\right)=\left(\begin{array}[]{l}\alpha_{0}z_{0}\\ \alpha_{1}z_{1}\\ \alpha_{2}z_{2}\\ \alpha_{3}z_{3}\\ \alpha_{4}z_{4}.\end{array}\right)$
In this case $L_{f}$ does not admit inequivalent Sasaki-Einstein metrics.
3.2. Rational homology 7-spheres: Thom-Sebastiani sums of invertible polynomials